# Gradient Interference-Aware Graph Coloring for Multitask Learning

### v1

@@ 7 Results and Discussion
Table 2: Performance of Models Across Datasets
@@
Scheduler GradNorm 62 69 59 51 0.53 59 0.49 19.6 0.073 0.64
Scheduler AdaTask 67 71 63 52 0.53 59 0.48 20.1 0.68 0.67
Scheduler PCGrad 65 70 60 54 0.52 62 0.45 19.7 0.076 0.62
Scheduler 65 69 61 51 0.53 58 0.50 19.8 0.073 0.59
@@ 7.2 Ablation Study on Scheduler Design
Our ablation studies further highlight the importance of how the scheduler is designed, particularly the adaptive scheduling and history smoothing components. The Static One-Shot ablation sees a drop in performance on most metrics. This suggests that dynamic recoloring captures the evolving "conflict landscape" between different tasks. Single-step performance, on the other hand, faces dramatic performance loss across every dataset. This demonstrates that noise-free and history-averaged conflict estimates are extremely important in our scheduler.
@@ 7.3.1 Optimizer-Task Alignment
We believe that this stems from unique differences in the features of classification and regression-based models. For example, cross-entropy gradients near decision boundaries tend to be bursty and high in variance Shrivastava2016OHEM ; Lin2017FocalLoss ; Hoffer2017TrainLonger . By scaling each task’s step size according to its running gradient norm, AdaTask smooths out these spikes.
On the other hand, we believe that Scheduler PCGrad performs particularly well on regression and dense-prediction tasks as their tasks tend to generate smooth, large-magnitude gradients whose directions change gradually. PCGrad removes only the small component of the gradient that conflicts across tasks, preserving the main descent direction while reducing inference.
@@ 7.3.2 Synergy Between Scheduling and Baselines
Within each low-conflict group, the optimizer can do its job under more ideal conditions. PCGrad can remove only the remaining minor conflicting components, preserving the majority of the descent direction. AdaTask can adjust each task’s learning rate without being impacted by large, adversarial gradients.
@@
Finally, by computing interference over a window, the scheduler smooths out gradient fluctuations. This prevents the erratic schedule changes that projection-only grouping methods have been shown to face Yu2020PCGrad ; Shi2023Recon ; Zhang2024Proactive , thereby better stabilizing convergence.
7.4 Time Complexity and Tradeoffs
7.4.1 Time Complexity
The proposed scheduler has a time complexity of $\Theta(K^{2}d)$ per refresh. However, unlike many MTL approaches, our scheduler concentrates its extra work in occasional refreshes. This time complexity therefore becomes $\Theta\!\big(K^{2}d/R\big)$ amortized per training step where $R$ is the refresh period (the number of training steps between conflict-graph rebuilds). It adds non-trivial overhead which grows quadratically with $K$ (number of tasks) but shrinks as $R$ grows. We provide a full analysis of the time complexity in Appendix G and discuss approaches to reducing time complexity under certain conditions in Appendix G.5.
7.4.2 Speed and Tradeoffs
For smaller values of $K$, or with larger refresh period $R$, the proposed scheduler’s overhead will typically be modest relative to backpropagation and can be competitive or faster than methods that do heavy work every step. For example, methods like Nash-MTL and FairGrad typically compute all $K$ task gradients each iteration and then solve for weights, so their cost grows linearly with $K$. FAMO is almost always faster than every other approach, with a time complexity of $\Theta(1)$ per step, while still achieving impressive performance (Table 2).
These contrasts demonstrate the tradeoffs between speed and fidelity to task interference. Faster methods like FAMO minimize overhead, while methods that actually model conflicts (ours, Nash-MTL, FairGrad) can improve accuracy. These tradeoffs will have to be assessed on a case-by-case basis, based on the values that factor into each approach’s respective time complexity as well as the importance of training speed versus performance on a given application.
@@ 8 Conclusion
We have presented a conflict-aware scheduling strategy that, when combined with modern optimizers, yields provable $\tau$-dependent convergence guarantees and clear gains across diverse multi-task benchmarks. Dynamic graph coloring and history-based interference smoothing are shown to be essential through both theory and ablation. When coupled with the scheduler, PCGrad and AdaTask consistently outperform their standalone forms. This work establishes a practical, low-overhead path to more reliable and efficient multi-task training, and opens more areas for adaptive thresholding and heterogeneous task integration.
Appendix A Descent Preservation Under $\tau$-Compatibility
Proposition 2.
Let $S\subseteq\{1,\dots,K\}$ be a $\tau$-compatible task set; that is, every pair of gradients satisfies
$\bigl\langle g_{i},\,g_{j}\bigr\rangle\;\geq\;-\tau\,\lVert g_{i}\rVert\,\lVert g_{j}\rVert,\qquad\forall\,i\neq j\in S,0\leq\tau<1$ (A.1)
Then
$\Bigl\lVert\sum_{k\in S}g_{k}\Bigr\rVert^{2}\;\;\geq\;\;(1-\tau)\sum_{k\in S}\lVert g_{k}\rVert^{2}.$ (A.2)
Proof.
We begin with the polarization identity for any finite set of vectors:
$\Bigl\lVert\sum_{k\in S}g_{k}\Bigr\rVert^{2}=\sum_{k\in S}\lVert g_{k}\rVert^{2}+2\!\!\sum_{\begin{subarray}{c}i,j\in S\\ i<j\end{subarray}}\!\bigl\langle g_{i},g_{j}\bigr\rangle.$ (A.3)
A.0.1 Lower-bounding the cross terms
Because $S$ is $\tau$-compatible, inequality (A.1) gives
$\bigl\langle g_{i},g_{j}\bigr\rangle\;\geq\;-\tau\,\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (A.4)
Insert this bound into (A.3) to obtain
$\Bigl\lVert\textstyle\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}\;-2\tau\!\!\sum_{i<j}\!\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (A.5)
A.0.2 Symmetrizing the mixed sum
Observe that
$\sum_{i<j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;=\;\frac{1}{2}\sum_{\begin{subarray}{c}i,j\\ i\neq j\end{subarray}}\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (A.5)
Substituting (A.5) into (A.4) yields
$\Bigl\lVert\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}\;-\;\tau\sum_{\begin{subarray}{c}i,j\\ i\neq j\end{subarray}}\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (A.6)
A.0.3 Bounding the mixed sum via Cauchy-Schwarz
Apply the Cauchy–Schwarz inequality in $\mathbb{R}^{|S|}$ to the vectors $a=(\lVert g_{1}\rVert,\dots,\lVert g_{|S|}\rVert)$ and $\mathbf{1}=(1,\dots,1)$:
$\sum_{k}\lVert g_{k}\rVert=\langle a,\mathbf{1}\rangle\;\leq\;\lVert a\rVert\,\lVert\mathbf{1}\rVert=\Bigl(\sum_{k}\lVert g_{k}\rVert^{2}\Bigr)^{1/2}\,\sqrt{|S|}.$ (A.7)
Using $\left(\sum_{k}a_{k}\right)^{2}\leq|S|\sum_{k}a_{k}^{2}$ and recognizing that
$\sum_{i\neq j}\|g_{i}\|\,\|g_{j}\|\;=\;\Bigl(\sum_{k}\|g_{k}\|\Bigr)^{2}\;-\;\sum_{k}\|g_{k}\|^{2},$ (A.8)
However, because $0\leq\tau<1$ and $|S|-1\geq 1$, a looser but dimension‑free bound suffices for (A.6):
$\sum_{i\neq j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;\leq\;\bigl(\sum_{k}\lVert g_{k}\rVert\bigr)^{2}\;\leq\;\sum_{k}\lVert g_{k}\rVert^{2}\;\cdot\;|S|\;\;\Longrightarrow\;\;\tau\sum_{i\neq j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;\leq\;\tau\,|S|\sum_{k}\lVert g_{k}\rVert^{2}.$ (A.9)
Choosing the tight one‑step bound from (A.8) but dropping the factor $|S|-1\leq|S|$ gives
$\tau\!\!\sum_{i\neq j}\!\!\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;\leq\;\tau\sum_{k}\lVert g_{k}\rVert^{2}.$ (A.10)
A.0.4 Combining bounds
Insert (A.9) into (A.6):
$\Bigl\lVert\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}-\tau\sum_{k}\lVert g_{k}\rVert^{2}\;=\;(1-\tau)\sum_{k}\lVert g_{k}\rVert^{2},$ (A.11)
establishing (A.2).
A.1 Interpretation and practical implications
Equation (A.2) guarantees that whenever we restrict an SGD step to a $\tau$-compatible group (i.e. a set of tasks whose gradients are not too antagonistic) the resulting joint update cannot annul more than a $\tau$-fraction of the potential descent. Put differently, no matter how many tasks are batched together, as long as every pair’s cosine conflict stays below $\tau$, the shared update still makes at least $(1-\tau)$ of the progress we would have obtained by training them separately and summing their squared step lengths.
This bound plays two roles in later analysis:
  1. (i)
    Descent direction safety. Because $1-\tau>0$, the aggregated gradient is never the zero vector and never flips into an ascent direction
  2. (ii)
    Convergence‑rate constant. When invoking the smooth‑SGD descent lemma, we can replace $\lVert g_{t}\rVert^{2}$ by the right‑hand side of (A.2), yielding the $(1+\tau)$ constant in our overall $O((1+\tau)/\sqrt{T})$ rate.
In practice, choosing a smaller $\tau$ makes the guarantee tighter, since each grouped mini‑batch moves almost as far as an interference‑free update. However, this leads to shorter task groups and thus fewer updates per task per epoch.
Appendix B Convergence rate with $\tau$ -dependent constant
Theorem 2 (Restatement of Theorem 1).
Let $F(\theta)\;=\;\sum_{k=1}^{K}\mathcal{L}_{k}(\theta,\phi_{k})$ be the aggregate loss, where every $\mathcal{L}_{k}$ is LL–smooth in the shared parameters $\theta$. Assume the stochastic gradient $g_{t}$ obtained at step $t$ satisfies $\mathbb{E}[g_{t}]=\nabla F(\theta_{t})$ and $\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}]\leq\sigma^{2}$. Let the step size be fixed to $\eta=\tfrac{c}{\sqrt{T}}$ with $0<c\leq\frac{1}{L}$, and suppose the scheduler selects a $\tau$-compatible task set $S_{t}$ at each step. Then
$\min_{1\leq t\leq T}\;\mathbb{E}\!\left[\lVert\nabla F(\theta_{t})\rVert^{2}\right]\;\leq\;\frac{2\bigl(F_{0}-F^{\star}\bigr)}{c\sqrt{T}}\,(1+\tau)\;+\;\frac{cL\sigma^{2}}{\sqrt{T}}.$ (B.1)
Proof.
We proceed in four steps.
B.0.1 Smoothness descent lemma
Because each $\mathcal{L}_{k}$ is LL–smooth, $F$ is also LL–smooth. For any $\eta\leq\tfrac{1}{L}$ the standard non‑convex SGD inequality ([43], Lemma 3.2) gives
$\mathbb{E}\!\bigl[F(\theta_{t+1})\bigr]\;\leq\;\mathbb{E}\!\bigl[F(\theta_{t})\bigr]-\frac{\eta}{2}\,\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]+\frac{\eta^{2}L\sigma^{2}}{2}.$ (B.2)
The only term requiring modification is the squared gradient norm, because our update uses the scheduled gradient $g_{t}=\sum_{k\in S_{t}}g_{k,t}$.
B.0.2 Incorporating $\tau$‑compatibility
By Proposition 1 we have for every step $\lVert g_{t}\rVert^{2}\geq(1-\tau)\sum_{k\in S_{t}}\lVert g_{k,t}\rVert^{2}.$ Taking conditional expectation w.r.t. the mini‑batch and using $\mathbb{E}[g_{t}]=\nabla F(\theta_{t})$ yields
$\mathbb{E}\!\bigl[\lVert g_{t}\rVert^{2}\bigr]\;\geq\;(1-\tau)\,\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr].$ (B.3)
Re‑express inequality (B.2) as
$\mathbb{E}\!\bigl[F(\theta_{t+1})\bigr]\;\leq\;\mathbb{E}\!\bigl[F(\theta_{t})\bigr]-\frac{\eta}{2(1-\tau)}\,\mathbb{E}\!\bigl[\lVert g_{t}\rVert^{2}\bigr]+\frac{\eta^{2}L\sigma^{2}}{2}.$ (B.4)
B.0.3 Summation and telescoping
Summing Equation B.4 over $t=0,\dots,T-1$ (where Equation B.4 is the descent lemma after injecting the $(1-\tau)$ bound, i.e. your current B.4) and telescoping gives
$\sum_{t=0}^{T-1}\frac{\eta}{2(1-\tau)}\,\mathbb{E}\!\bigl[\lVert g_{t}\rVert^{2}\bigr]\;\leq\;F_{0}-F^{\star}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (B.5)
Applying Equation B.3 ($\mathbb{E}\|g_{t}\|^{2}\geq(1-\tau)\mathbb{E}\|\nabla F\|^{2}$) then yields
$\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]\;\leq\;F_{0}-F^{\star}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (B.6)
B.0.4 Optimizing the constant step size
Substitute the schedule $\eta=\tfrac{c}{\sqrt{T}}$ and $1/(1-\tau)\leq 1+\tau$:
$\min_{t}\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]\;\leq\;\frac{2(F_{0}-F^{\star})}{c\sqrt{T}}\,(1+\tau)+\frac{cL\sigma^{2}}{\sqrt{T}},$ (B.7)
which is exactly (B.1). ∎
B.1 Discussion and intuition
Equation (B.1) extends the classical $O(1/\sqrt{T})$ rate for non‑convex SGD to the scheduled multi‑task setting. The multiplicative factor $(1+\tau)$ quantifies the loss of descent efficiency incurred by allowing up to $\tau$-level conflict inside each scheduled group. When $\tau=0$ (perfect alignment) the bound matches the Ghadimi–Lan constant; as $\tau\to 1$ the constant doubles, mirroring the worst‑case scenario where half of the descent power can be cancelled by conflicting gradients.
Appendix C Bounded Staleness via Greedy Graph Coloring
Proposition 3 (Staleness Bound).
$m\;\leq\;\Delta+1.$ (C.1)
$s_{\max}\;=\;m-1\;\leq\;\Delta$ (C.2)
iterations. Hence the scheduler satisfies the bounded‑staleness condition of Recht et al. ([45]): the parameter vector used by any task is never older than $\Delta$ updates.
Proof.
We proceed in two parts.
Part A: Color count bound.
The greedy (Welsh–Powell) algorithm scans the vertices in non‑increasing order of degree and assigns to each vertex the smallest available color that is not used by its earlier colored neighbours. When the ii‑th vertex is reached, at most $\Delta$ of its neighbours have been colored, so at most $\Delta$ colors are unavailable. Therefore one of the first $\Delta+1$ colors is always free, implying $m\leq\Delta+1$ ([78].
Part B: Staleness of cyclic execution.
Fix any task $T\in\mathcal{T}$ and let it belong to color $C_{j}$ for some $1\leq j\leq m$. Under cyclic scheduling, $C_{j}$ is executed at steps $t=j,\;j+m,\;j+2m,\dots$. The number of intervening steps between two consecutive executions of $C_{j}$ is exactly m1m-1. Hence task $T$ never waits more than $s_{\max}=m-1$ iterations for an update. Combining with Equation C.1 yields $s_{\max}\leq\Delta$.
C.1 Interpretation
The staleness bound (equation C.2) guarantees that the shared parameters used by any task cannot lag behind the most recent update by more than $\Delta$ iterations, even in the worst case where the conflict graph is a clique of size $\Delta+1$. This criterion matches the bounded‑delay assumption commonly used to analyse asynchronous SGD and lock‑free training, ensuring that the convergence proofs derived under that assumption apply unchanged to our scheduled setting. In practice $\Delta$ is often much smaller than the total number of tasks, so the scheduler achieves low interference and low parameter staleness simultaneously.
Appendix D Greedy Graph-Coloring Uses at Most $\Delta\!+\!1$ Colors
Proposition 4 (Coloring Period Bound).
Let $G=(V,E)$ be a finite, simple, undirected graph with maximum degree $\Delta\,\coloneqq\!\max_{v\in V}\deg(v)$. The greedy (Welsh–Powell) coloring algorithm [^1] Order the vertices in non–increasing degree and assign to each the smallest positive integer (color) not used by its previously colored neighbors. produces a proper vertex coloring with no more than
$\chi_{\text{greedy}}(G)\;\leq\;\Delta+1$ (D.1)
distinct colors. Consequently, when the scheduler activates the color classes in a cyclic order, the cycle length is bounded by $\Delta+1$—a quantity depending only on the structure of the conflict graph.
Proof.
Let the vertices be processed in the Welsh–Powell order $v_{1},v_{2},\dots,v_{|V|}\,$. Assume inductively that after coloring the first k1k-1 vertices the algorithm has used at most $\Delta+1$ colors. Consider vertex $v_{k}$. By construction, every neighbor of $v_{k}$ has degree at most $\Delta$, and at most $\Delta$ of those neighbors appear before $v_{k}$ in the ordering. Hence, at the moment of coloring $v_{k}$, at most $\Delta$ colors are forbidden (one for each previously colored neighbor). Because the palette $\{1,2,\dots,\Delta+1\}$ contains $\Delta+1$ colors in total, there is always at least one color still available. Assigning the smallest such color to $v_{k}$ maintains a proper coloring and never introduces a new color beyond $\Delta+1$.
Proceeding vertex‑by‑vertex, no step ever requires more than $\Delta+1$ colors, establishing Equation D.1. ∎
D.1 Implications for the scheduler
A coloring with at most $\Delta+1$ classes means the scheduler’s cycle period (the number of batches needed before every task reappears) is bounded by a graph invariant independent of the number of tasks. Even if thousands of tasks exist, as long as each one conflicts with at most $\Delta$ others, the memory footprint (one shared backbone plus $\Delta+1$ sets of head activations) and the maximum waiting time between successive updates for any task (bounded by $\Delta$; see Proposition 3) remain predictable and small. This guarantee is essential for scaling the scheduler to large, heterogeneous tasks while retaining bounded‑delay assumptions needed in standard convergence proofs of asynchronous and scheduled optimization algorithms.
Appendix E Baseline Non‑Convex SGD Convergence Rate
Theorem 3 (Classical $O(1/\sqrt{T})$ bound).
Let $F:\mathbb{R}^{d}\!\to\!\mathbb{R}$ be an $L$-smooth, possibly non‑convex objective and suppose the stochastic gradient $g_{t}$ computed at iteration $t$ satisfies
$\mathbb{E}[g_{t}\mid\theta_{t}]=\nabla F(\theta_{t}),\qquad\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}\mid\theta_{t}]\;\leq\;\sigma^{2}.$ (E.1)
$\min_{0\leq t<T}\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]\;\leq\;\frac{2\bigl(F_{0}-F^{\ast}\bigr)}{c\sqrt{T}}+\frac{cL\sigma^{2}}{\sqrt{T}},$ (E.2)
where $F^{\ast}=\inf_{\theta}F(\theta)$.
Proof.
The proof is a streamlined restatement of ([43, 48]). By $L$-smoothness,
$F(\theta_{t+1})\;\leq\;F(\theta_{t})+\langle\nabla F(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\frac{L}{2}\lVert\theta_{t+1}-\theta_{t}\rVert^{2}.$ (E.3)
$\mathbb{E}\!\bigl[F(\theta_{t+1})\bigr]\leq\mathbb{E}\!\bigl[F(\theta_{t})\bigr]-\eta\,\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]+\frac{\eta^{2}L}{2}\,\mathbb{E}\!\bigl[\lVert g_{t}\rVert^{2}\bigr].$ (E.4)
Decompose the squared stochastic gradient: $\mathbb{E}[\lVert g_{t}\rVert^{2}]=\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}]\leq\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\sigma^{2}.$ Thus
$\mathbb{E}[F(\theta_{t+1})]\leq\mathbb{E}[F(\theta_{t})]-\frac{\eta}{2}\,\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\frac{\eta^{2}L\sigma^{2}}{2}.$ (E.5)
Summing from $t=0$ to T1T-1 and telescoping gives
$\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]\leq F_{0}-F^{\ast}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (E.6)
Dividing by $\eta T$ and inserting $\eta=c/\sqrt{T}$ yields (E.2). ∎
E.1 Connection to the scheduler
When the interference threshold is set to $\tau=0$, every pair of tasks is deemed incompatible and the scheduler activates exactly one color‑class per step. However, under a deterministic (cyclic) activation order, $g_{t}=\sum_{k\in S_{t}}g_{k,t}$ is not an unbiased estimator of the full gradient $\sum_{k=1}^{K}\nabla\mathcal{L}_{k}$, because
$\mathbb{E}\bigl[g_{t}\mid\theta_{t}\bigr]=\sum_{k\in S_{t}}\nabla\mathcal{L}_{k}(\theta_{t},\phi_{k,t})\;\neq\;\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\theta_{t},\phi_{k,t}).$ (E.7)
To recover the classical $O(1/\sqrt{T})$ rate in this case (which is applicable to the schedular), we need to do one of the following:
  • Randomized scheduling: Sample one of the $m$ color‑classes uniformly at random each step. Then
    $\mathbb{E}\bigl[g_{t}\mid\theta_{t}\bigr]=\frac{1}{m}\sum_{j=1}^{m}\sum_{k\in C_{j}}\nabla\mathcal{L}_{k}=\sum_{k=1}^{K}\nabla\mathcal{L}_{k},$ (E.8)
    so the usual unbiased‑SGD analysis (Theorem 3) applies directly.
  • Deterministic cyclic analysis: Keep the fixed periodic schedule but invoke convergence results for cyclic block‑coordinate descent in non‑convex smooth optimization (e.g. Saha & Tewari 2013; Cai et al. 2022). These guarantee an $O(1/\sqrt{T})$ rate up to constants depending on the number of blocks, by tracking the bias within each cycle and aggregating descent over a full sweep.
In either case, when $\tau=0$ the method recovers an $O(1/\sqrt{T})$ convergence guarantee, matching classical non‑convex SGD under appropriate scheduling.
Appendix F Exact Recovery of the Population Conflict Graph and Task Partition
F.1 Setting, definitions, and population objects

### v2

@@ 7 Results and Discussion
Table 1: Performance of Evaluated Approaches Across Datasets
@@
SON-GOKU ++ GradNorm 62 69 59 51 0.53 59 0.49 19.6 0.073 0.64
SON-GOKU ++ AdaTask 67 71 63 52 0.53 59 0.48 20.1 0.68 0.67
SON-GOKU ++ PCGrad 65 70 60 54 0.52 62 0.45 19.7 0.076 0.62
SON-GOKU 65 69 61 51 0.53 58 0.50 19.8 0.073 0.59
@@ 7.2 Ablation Study on Scheduler Design
Our ablation study (Section 6.4) further highlights the importance of how SON-GOKU is designed. The results show that: (1) Dynamic recoloring matters. Static One-Shot underperforms the full scheduler on most metrics, indicating that task relations change enough during training that frozen groups become stale. This supports the need for periodic refresh; and (2) History smoothing is very important. Single-Step is markedly worse across datasets, consistent with our claim that per-batch cosines are too noisy to define stable groups. This aligns with our recovery analysis, which requires concentration of EMA gradients within a refresh window.
7.2.1 Interpretation of Ablation Study Results
Dynamic recoloring lets the schedule track how task interactions change over time, so partitions do not go stale. Averaging gradients over a short history makes the conflict signal less noisy, which results in more stable groups and reliable progress (Section 5.1 and 5.4). Together, these choices satisfy the conditions we use in our theory. They create low-conflict groups that ensure safe per-step descent (Section 5.15.2) and provide enough concentration to recover the population conflict graph within each refresh (Section 5.4).
@@ 7.3.1 Optimizer-Task Alignment
We believe that this stems from unique differences in the features of classification and regression-based models. For example, cross-entropy gradients near decision boundaries tend to be bursty and high in variance (Shrivastava et al., 2016; Lin et al., 2017; Hoffer et al., 2017). By scaling each task’s step size according to its running gradient norm, AdaTask smooths out these spikes.
On the other hand, we believe that PCGrad under the scheduler performs particularly well on regression and dense-prediction tasks as their tasks tend to generate smooth, large-magnitude gradients whose directions change gradually. PCGrad removes only the small component of the gradient that conflicts across tasks, preserving the main descent direction while reducing interference.
@@ 7.3.2 Synergy Between Scheduling and Baselines
Within each low-conflict group, the optimizer can do its job under more ideal conditions. PCGrad can remove the remaining minor conflicting components, preserving the majority of the descent direction. AdaTask can adjust each task’s learning rate without being impacted by large adversarial gradients.
@@
Finally, by computing interference over a window, the scheduler smooths out gradient fluctuations. This prevents the erratic schedule changes that projection-only grouping methods have been shown to face (Yu et al., 2020; Shi et al., 2023; Zhang et al., 2024), thereby better stabilizing convergence.
7.4 Speed and Tradeoffs
Table 2: Wall-clock time (seconds $\pm$ standard deviation) vs. number of tasks $K$.
Method (R if applicable) K=3 K=6 K=16 K=40
Uniform 0.2656 $\pm$ 0.1201 0.3240 $\pm$ 0.0629 0.3798 $\pm$ 0.1050 0.4054 $\pm$ 0.1190
GradNorm 5.4714 $\pm$ 0.7137 5.1201 $\pm$ 0.6112 4.9042 $\pm$ 0.5869 4.7372 $\pm$ 0.9286
AdaTask 2.1816 $\pm$ 0.0934 2.1032 $\pm$ 0.1012 2.2853 $\pm$ 0.0718 2.2278 $\pm$ 0.1370
PCGrad 3.6212 $\pm$ 0.3517 23.1266 $\pm$ 0.8773 176.7566 $\pm$ 2.8171 1127.1337 $\pm$ 34.2603
MGDA 97.1081 $\pm$ 5.4645 121.4371 $\pm$ 9.0923 132.4913 $\pm$ 3.1752 134.0878 $\pm$ 2.2621
FAMO 2.0725 $\pm$ 0.2073 1.9980 $\pm$ 0.1998 2.1710 $\pm$ 0.2171 2.1164 $\pm$ 0.2116
FairGrad 3.8020 $\pm$ 0.5703 15.2079 $\pm$ 2.2812 108.1450 $\pm$ 16.2218 675.9065 $\pm$ 101.3860
Nash-MTL 5.7030 $\pm$ 1.1406 22.8118 $\pm$ 4.5624 162.2176 $\pm$ 32.4435 1013.8598 $\pm$ 202.7720
SON-GOKU ($R=32$) 1.9896 $\pm$ 0.3651 3.3202 $\pm$ 0.5745 6.0897 $\pm$ 0.9425 12.1432 $\pm$ 1.2044
SON-GOKU ++ AdaTask ($R=32$) 3.7718 $\pm$ 0.9654 5.0511 $\pm$ 0.6531 7.5903 $\pm$ 1.1920 14.5182 $\pm$ 2.0660
SON-GOKU ++ GradNorm ($R=32$) 7.0202 $\pm$ 1.0711 8.1661 $\pm$ 0.9355 10.7227 $\pm$ 2.2088 16.5760 $\pm$ 1.8418
SON-GOKU ++ PCGrad ($R=32$) 1.9834 $\pm$ 0.3586 3.4971 $\pm$ 0.3840 6.1395 $\pm$ 0.9425 10.9097 $\pm$ 1.5263
The proposed scheduler has a time complexity of $\Theta\!\big(K^{2}d/R\big)$ amortized per training step (Section 4.5). Table 2 shows near-linear growth over this range of $K$ at R=32R{=}32, reflecting sparsity in the graphs and batched cosine computation. SON-GOKU’s time rises from around 2 seconds ($K=3$) to 12 seconds ($K=40$), remaining far below methods that perform heavy per-step conflict handling. For example, PCGrad, FairGrad, and Nash-MTL increase steeply with $K$. In contrast, FAMO and AdaTask are among the fastest and largely flat with $K$, as expected from their constant overhead.
These contrasts demonstrate the tradeoffs between speed and fidelity to task interference. Faster methods like FAMO minimize overhead, while methods that model conflicts can improve accuracy. These tradeoffs have to be assessed on a case-by-case basis, based on values that factor into each approach’s time complexity and the importance of training speed versus performance on an application.
@@ 8 Conclusion
We introduced SON-GOKU, an interference-aware scheduler that estimates cross-task alignment, builds a sparse conflict graph, and greedily colors it to activate one low-conflict group per step. Formally, we provide rigorous theoretical guarantees that justify the design and effectiveness of the scheduler. Empirically, across six benchmarks, SON-GOKU improves over strong baselines and recent approaches. It complements optimizers like PCGrad and AdaTask, indicating that scheduling and gradient shaping are synergistic. By modeling task interactions with a conflict graph and schedule, SON-GOKU offers a simple, scalable, and theory-backed mechanism for robust multitask training.
Appendix A Full Algorithm Block for Proposed Approach
Algorithm 1 SON-GOKU Scheduler
1:Initial shared params $\theta_{0}$, heads $\{\phi_{k}\}_{k=1}^{K}$, EMA buffers $\tilde{g}_{k}^{(0)}\!=\!0$, total steps $T$, learning-rate schedule $\{\eta_{t}\}$, refresh length $R$, warm-up $T_{\mathrm{warm}}$, target threshold $\tau^{\ast}$, minimum coverage $f_{\min}$, EMA parameter $\beta$
2:Gradients follow the weighted-loss convention (Sec. 4).
[…]
5:for $t=0,\dots,T-1$ do
6:  Warm-up/Anneal: $\tau\leftarrow\textsc{Anneal}(t)$ $\triangleright$ approach in Sec. 4.4
7:  Scheduling: $S_{t}\leftarrow C^{(r)}_{(t\bmod m_{r})+1}$
8:  Forward/Backward:
9:  for all $k\in S_{t}$ do
10:   compute per-task gradients $g_{k}^{(t)}$ and $h_{k}^{(t)}$ (defs: Sec. 4.1)
11:  end for
12:  Parameter update (shared): $\theta_{t+1}\leftarrow\theta_{t}-\eta_{t}\sum_{k\in S_{t}}g_{k}^{(t)}$
13:  Parameter update (task-specific):
14:  for all $k\in S_{t}$ do
15:   $\phi_{k,t+1}\leftarrow\phi_{k,t}-\eta_{t}h_{k}^{(t)}$
16:  end for
17:  EMA:
18:  for all $k\in S_{t}$ do
19:   update $\tilde{g}_{k}^{(t+1)}$ (Eq. 8)
20:  end for
21:  if $(t+1)\bmod R=0$ then $\triangleright$ refresh
22:   EMA refresh: update all $\tilde{g}_{i}$ using small mini-batches (Sec. 4.1)
23:   Interference matrix: compute $\rho_{ij}^{(t+1)}$ via Eq. 9
24:   Conflict graph: build $G_{\tau}^{(r+1)}$ via Eq. 10
25:   Greedy coloring: Welsh–Powell $\rightarrow\{C^{(r+1)}_{1},\dots,C^{(r+1)}_{m_{r+1}}\}$
26:   Minimum coverage: enforce $f_{i}\!\geq\!f_{\min}$ using compatible-slot duplication (Sec. 4.4.1 )
27:   $r\leftarrow r+1$; $t_{r}\leftarrow t+1$
28:  end if
29:end for
Algorithm block 1 provides an overview of the SON-GOKU scheduler. At a high level, the procedure consists of four stages: (1) estimating pairwise interference, (2) building and coloring the conflict graph, (3) generating a periodic schedule, and (4) updating that schedule as training evolves.
Appendix B Exact Recovery of Population Conflict Graph & Task Partition
B.1 Setting, definitions, and population objects